OpenStudy (anonymous):

Theo is tracking the growth of his plant. When he got the plant, it was 2 cm high. It grows 0.5 cm per day. Which graph represents the growth of the plant?

OpenStudy (anonymous):

f(x) = 0.5*x + 2, where x is the number of days and f(x) is the high of the plant in day x

OpenStudy (anonymous):

for example, at day x=0, the plant will be f(0) = 2 cm at day x=1, f(1) = 0.5+2 = 2.5 cm